# Towards Browser Controls to Protect Cookies from Malicious Extensions

### v2

@@ 7.3 Server-Side
The BrowserOnly attribute defines how cookies are accessed. Thus BrowserOnly cookies require no additional processing on the server. Monitored cookies, on the other hand, require changes to how servers create and parse cookies. When creating a Monitored cookie, the server must encrypt the Monitored message. Using AES-256-CBC, we found the average time to encrypt to be 6.7 $\mu$s. Monitored cookies also introduce new server-side checks to ensure the report and cookie are valid. While the report verification is always the same, cookie validation depends on the server’s policy. In our experiments, we checked that the changelog was empty and the cookies settings matched an expected configuration. These checks added 3,419 $\mu$s to the server-side processing time.
7.4 Latency
@@
We also measure the time taken to build requests and parse responses with different cookies. For this, we timed how long it took the URLRequestHttpJob to handle the message as only this portion of the network stack changed. Timing this portion alone also helps reduce noise caused by the rest of the browser. Unlike the previous tests, timing the network stack’s behavior required reloading the test page for each sample. As such, we only gathered 10,000 samples to calculate the median in these experiments (depicted in Figure 5(A)).
@@
Parsing responses in CREAM requires BrowserOnly cookies be stripped from responses before any extensions run. Due to this, parsing responses with a BrowserOnly cookie took 133 $\mu$s (4.5%) longer than the baseline. Parsing responses with either a standard or a Monitored cookie has no new behavior. As such, parsing either was on par with the baseline, 14 and 21 $\mu$s faster respectively. Similar to the individual cookie operations, this improvement is due to variability in browser processes. When building the request BroswerOnly cookies must be withheld until after extensions run, Monitored reports must be generated, and the browser key must be added to the headers (if an HTTPS request). When using standard cookies only, CREAM was 55 $\mu$s (3.2%) slower than the baseline despite the only new behavior being attaching the key which took 8 $\mu$s on average. Using BrowserOnly and Monitored cookies took 58 and 220 $\mu$s (3.4% and 12.8%) longer than the baseline. Monitored cookies additional overhead stems from the report generations. As such, we also timed the report generation itself (Shown in Figure 5(B)). Generating a report with no changes took 47 $\mu$s on average. Reports with 1 and 10 changelog entries took 49 and 62 $\mu$s respectively. Finally generating a report with 5 Monitored cookies took 107 $\mu$s on average. This increase comes from running the sub-report generation process multiple times.
8 Privacy Considerations
@@
CREAM’s design aims to strike a good balance between security and privacy. Below we discuss privacy-related considerations.
@@
Private/Incognito Browsing. To limit tracking cookies [76, 77, 78], browsers restrict which contexts cookies can be accessed from. While we discuss the Cookie Monster as a single entity, browsers have different Cookie Monsters for different browsing contexts. For example, Incognito browsing has a separate Cookie Monster from standard browsing. Thus, Incognito cookies are not accessible during standard browsing (and vice-versa). As discussed in Section 2.3, cookies can also be partitioned to restrict their accessibility by embedding context. As CREAM works on individual cookies, it does not inhibit and remains compatible with these privacy features. Further, CREAM-protected cookies remain manageable by users via the normal browser management interfaces.
@@
Privacy-Enhancing Extensions. Some extensions seek to protect user privacy. Among other checks, they often use the Cookies API to search for and remove known tracking cookies from the browser/requests. As the BrowserOnly attribute hides cookies from extensions, it could be leveraged by servers to avoid privacy-enhancing extensions. As a middle ground, it is possible to allow users to selectively disable the BrowserOnly attribute for specific trusted extensions. This opt-out approach would protect the user by default while allowing discretionary disablement. That said, we note that, even without the BrowserOnly attribute, user tracking is still possible through techniques such as browser fingerprinting [79, 80] which are orthogonal to our goal. Further, without approaches to protect cookies, session cookies can be stolen/used, leading to accounts (and all sensitive information therein) being compromised, i.e., a very strong privacy violation. Without browser-based controls such as CREAM, the latter is possible irrespective of privacy-enhancing extensions.
@@
Tracking and Fingerprinting. Monitored cookies rely on cryptographic keys to tie a cookie to a browser. However, as keys are generated per site, they cannot be used to track the user across different domains. It is also possible to generate keys per origin to limit tracking within a single site, e.g., similar to browsers’ origin isolation [81]. We note that, as session cookies contain the user’s authenticated ID, they identify the user. Thus, irrespective of CREAM, while a single session cookie is only valid for a single session, multiple session cookies can still track the user across sessions. However, it is also possible to update/refresh the keys for each session to prevent their use for cross-session tracking. These keys could also allow individual sites to track users across browsing contexts. This can be prevented by generating keys per site and context rather than per site alone. However, this would incur additional runtime and storage costs.
9 Related Work
@@
Session Protection is a similar external control offered by Flask-Login [87] (Flask’s session management module). When enabled, the server hashes the user’s IP address and user agent (browser) for each session cookie. These hashes are stored on the server and link the cookies to specific hosts. Then Flask recalculates the hash for all requests with a session cookie and compares it against the saved hash. In the strictest configuration, a hash mismatch invalidates the session. However, by default, a mismatch only marks the session as stale. In this state, the user can still access the site but must re-authenticate for any action that requires a "fresh" session. Further, Flask also supports a separate "Remember Me" token to save and restore sessions. This token is not protected by Session Protection and thus can be stolen and used to bypass Session Protection altogether. Due to this, the Flask-Paranoid extension [88] extends Session Protection to the Remember Me token as well. Regardless, while Session Protection can detect cookie theft, it cannot accurately determine if a cookie was stolen as benign actions (i.e., IP address changes) also violate the control. Thus stolen cookies could still provide access to most of the session.
Other proposals use extensions to bolster the security of cookies [4, 89, 90]. CookieExt [4] detects session cookies in incoming requests and attaches the Secure and HttpOnly attributes to them. Serene [89] maintains a database of session cookie values and ensures the value of outbound cookies are in the database. As discussed in Section 1, malicious extensions can bypass existing cookie attributes and reattach cookies to requests. Thus the aforementioned extension-based approaches cannot protect against other extensions.
Many controls change cookies themselves [91, 2, 92, 93]. Signed cookies contain a cryptographic signature in the cookie’s value to support integrity-checking [92] and access-control [93]. Origin cookies [91] add a new “Origin" attribute that greatly limits a cookie’s accessibility to only the origin that set it. One-Time Cookies (OTC) [2] generate a unique token per request rather than per session. OTCs are then tied to a request with a MAC and verified with a shared session key. This session key and any other information needed for verification are stored in a session ticket, encrypted by the server’s secret key. These features limit how cookies are accessed, but do not protect against and still allow theft by malicious extensions. While OTCs reduce the usability of a stolen cookie to a single request, they are enabled through HTTP headers. Thus, malicious extensions can disable OTCs outright and force standard cookie authentication.
@@
10 Conclusion
We proposed CREAM: a modification to the web browser’s trust model to protect cookies from malicious extensions. CREAM introduces two new cookie attributes: BrowserOnly and Monitored. BrowserOnly prevents cookies from being accessed by extensions and JS outright. Monitored maintains accessibility but ties the cookie to a single browser, logs all changes made to the cookie, and sends an authenticated report of changes to the server alongside the cookie. This allows the server to detect and refuse stolen and tampered cookies. We provide an open-source prototype of CREAM built atop Chromium [39] and evaluate it to show that both the BrowserOnly and Monitored attributes incur little overhead.
@@ Appendix A Scraping Browser Extensions
To analyze Chrome’s and Firefox’s extensions we performed a series of iterative scrapes. We began by manually scraping both the Chrome Web Store and AMO, collecting each extension’s name, user count, and URL. Both extension repositories divide extensions into several overlapping categories. Thus, for the AMO, we filtered extensions by category and iterated over all matching results. Doing this, we pulled an initial set of 35,098 Firefox extensions. Scraping the Web Store proved more difficult as search results were loaded dynamically. This required continually pressing a button labeled "load more" to view more results. Further complicating the process, after 50 presses, the webpage would refresh and revert to the initial set of results. Due to this, it was impossible to scrape all Chrome extensions from the search results directly. Instead, we used JS to grab an initial list of 30,473 extension URLs. Then, while scraping the name and user count from each URL, we also pulled any URLs included in the “related extensions” portion of the page. Any previously unseen URLs were added to the scrape. When finished, we found 69,203 Chrome extensions in total.
Compared to prior work this number was low. Between 2018 and 2022, several studies stated the Web Store contained around 130,000 extensions [98, 99, 13] while Chrome themselves reported having over 180,000 [100]. Chrome’s larger estimate likely included both Themes and Chrome Apps. We discuss Themes and Chrome Apps in greater detail later in this section. Due to the discrepancy, we identified a prior scrape performed in 2021 containing 130,445 extension URLs [101] and compared our results against it. From this comparison, we discovered 90,152 previously unseen URLs, however only 20,877 of them still worked presently.
Despite increasing our total Chrome extensions to 90,080, this was still relatively low. Another study on privacy inconsistencies in extensions [98] used sitemaps to discover extensions. Using the sitemap in [102], we found 172,386 extensions. Interestingly, when comparing these results against our prior scrapes, 2,151 previously discovered extensions were missing. This suggests they had been removed from the Web Store between scrapes. We also found AMO’s sitemap [103] and re-scrapped it resulting in 533,768 extensions. Similar to Chrome, 1,170 previously identified extensions were not in the sitemap and were likely removed.
Table 4: Breakdown of manifest versioning, host permissions, and access to cookies
@@
A.1 Categorizing extensions
As discussed, we identified 172,386 Chrome and 533,768 Firefox extensions during our scrape. However, the majority of these extensions are "Themes". Themes are purely cosmetic and contain no JS or HTML [104]. For this reason, we exclude Themes from our analysis. Similarly, 16,380 Chrome extensions were unlabeled. We believe these to be Chrome Apps, a now deprecated subclass of extensions that behave similarly to native applications [105]. Though deprecated, Apps are still currently available on Enterprise and Education ChromeOS builds [106]. Hence, we do analyze the Chrome App’s permissions but separately from our main discussion. 22 Firefox extensions were labeled as Android. Given the distinction, we again analyze their permissions separately from the main discussion. Filtering out these categories we were left with 122,951 Chrome and 37,783 Firefox extensions.
A.2 Scraping Manifests
With our dataset of 215,941 extensions (including Chrome Apps and Android variants), the final step was to scrape each extension’s manifest file. Using an online extension viewer [107], we entered each extension’s URL, opened its files, and downloaded its manifest. Despite splitting the dataset across 10 threads, the initial execution took about 36 hours. We then repeated the process for all URLs which resulted in an error with more time between each action. These later runs helped remove false errors due to rate limiting and provided more time for larger extensions to load. When complete, we downloaded 119,152 Chrome and 37,631 Firefox extension manifests. Attempts to scrape the remaining manifests resulted in various 400 HTTP errors, suggesting they were private or removed from the repository.
@@ Appendix B Chrome Apps & Android Analysis
As discussed in Section 3.2, we discovered 16,380 Chrome Apps and 22 Firefox for Android extensions. From this, we pulled 12,156 App and 22 Android manifests and analyzed them for their version, host permissions, and to see if they have access to cookies. A summary of our findings can be found in Table 4.
Nearly all manifests analyzed (99.5%) were written using manifest V2. This is likely due to a lack of incentive to change formats as Firefox currently has no plans to deprecate the format and Chrome Apps will be fully deprecated in 2025 [105]. Looking at host permissions, only 4.3% of manifests declared broad host permissions. However, most Android extensions (81.8%) had “$<$all_urls$>$" permissions. Only 4.1% of Chrome Apps had any broad host permissions. This lack of permission may be due to Chrome Apps performing more like native applications thus running more locally rather than needing to access many websites.
We saw similar low numbers concerning access to cookies. Only 17 manifests (0.1%) had the Cookies permission. Similarly, only 41 (0.3%) used the WebRequest API with 26 (0.2%) also having access to the WebRequestBlocking permission. While low, 63.6% of Android extensions had the WebRequest permission, and all but one of these also declared the WebRequesBlocking permission. Despite this, manifests with the Cookies, WebRequest, and WebRequestBlocking API affected a combined 2,245,591, 14,699,240, and 13,622,263 users respectively. No manifest declared the DeclarativeNetRequest permission. We further analyzed which manifests that could access cookies also had “$<$all_urls$>$" host permissions. Table 5 summarizes these findings. Both Android extensions with the Cookies permission could access all cookies in the browser. This was less common in Chrome Apps with only 20% of those with Cookies permissions able to access all cookies. 92.9% of Android extensions and 48.1% of Chrome Apps with WebRequest permissions can see all traffic in the browser. Similarly, all of these Android extensions and 60% of these Chrome Apps also declare the WebRequestBlocking permission.
Table 5: Manifests with ”$<$all_urls$>$” and cookie access

### v3

@@ VII-C Server-Side
We also measure both attributes’ impact on server-side operations. The BrowserOnly attribute defines how cookies are accessed in the browser. Thus BrowserOnly cookies do not add any overhead to the server. Monitored cookies on the hand, change how servers create and parse cookies. When creating a Monitored cookie, the server must encrypt the corresponding Monitored message. Using AES-256-CBC, this encryption took 6.7 $\mu$s on average. Monitored cookies also add new server-side checks to validate the cookie and report. While the report’s MAC is always verified in the same way, cookie validation depends on the server’s policy. In our experiments, a cookie was valid if its changelog was empty and current settings matched an expected configuration. For a single user, these checks added 256 $\mu$s to the server-side processing time. To determine the impact of this additional functionality under load, we also measure the cookie validation under various numbers of concurrent users. Using the httperf tool [83], we tested the server with 10, 100, and 1000 concurrent users. In each successive test, the processing time scaled by a factor of 10 (2,600, 28,151, and 298,873 $\mu$s respectively). Thus the Monitored attributes additional server-side overhead scales linearly with the number of concurrent users.
Along with direct processing times, we measure the impact of each attribute on the server’s maximum throughput (shown in Figure 5.) Using standard cookies, the server’s throughput caps at 7,545 replies/s on average. Varying the number of concurrent connections from 1000 to 5000 had a negligible impact on the server’s throughput. BrowserOnly cookies achieved similar throughput in most tests. However, when testing BrowserOnly cookies with 5000 users, the throughput dropped to 6861 replies/s. We believe this decrease reflects variability on the test device itself rather than an actual drop in throughput as a similar decline was not seen between Monitored tests. For Monitored cookies, the server’s throughput capped at an average of 6,147 replies/s. This decrease is due to the additional server-side checks performed on Monitored cookies. We note that our throughput results represent a comfortable lower bound, as tests were performed (for the sake of comparison with the baseline) on a single-core experimental Node.js server. Naturally, throughput should increase with server-side resources and load distribution across multiple server-side cores/machines.
VII-D Latency
@@
Finally, we measure the time taken to build requests and parse responses with different cookies. For this, we timed the URLRequestHttpJob behavior as only this portion of the network stack changed. Timing this portion alone also helps reduce noise caused by the rest of the browser. Unlike the previous tests, timing the network stack’s behavior required reloading the test page for each sample. As such, we only gathered 10,000 samples to calculate the median in these experiments (depicted in Figure 6).
11111.51.51.51.522222.52.52.52.53333BaselineStandardBrowserOnlyMonitoredTime (ms)RequestResponse
Figure 6: Latency of parsing network traffic with different types of cookies
000.050.050.050.050.10.10.10.10.150.150.150.150.20.20.20.2Empty110100Time (ms)
(A) Chagelog Entries
000.50.50.50.511111.51.51.51.522222.52.52.52.511050100Time (ms)
(B) Cookies
Figure 7: Latency of generating Monitored reports with different numbers of (A) log entries and (B) cookies
@@
The only change to parsing responses in CREAM is that BrowserOnly cookies must be stripped from responses before extensions execute. As such, responses with a BrowserOnly cookie took 133 $\mu$s (4.5%) longer than the baseline to parse. Responses with standard or Monitored cookies require no new behavior. Thus parsing either was on par with the baseline, 14 and 21 $\mu$s faster respectively. Again this improvement is due to the variability in other browser processes. Building requests in CREAM requires withholding BrowserOnly cookies until after extensions execute, generating Monitored reports, and adding the browser’s key to the headers (HTTPS only). Only using standard cookies, CREAM was 55 $\mu$s (3.2%) slower than the baseline. Interestingly, attaching the browser’s key (the only new behavior in this case) only took 8 $\mu$s on average. Requests with BrowserOnly and Monitored cookies took 58 and 220 $\mu$s (3.4% and 12.8%) longer than the baseline respectively.
Monitored cookies’ additional overhead stems from generating the Monitored report. As such we also timed the report generation itself. Specifically, we measure the time to generate a report for a single cookie with various changlelog entries (shown in Figure 7(A)) and multiple cookies (shown in Figure 7(B)). Generating the report for a cookie with no changes took 50.8 $\mu$s. Reports with 1 and 10 changelog entries added little overhead taking 51.8 and 50.8 $\mu$s respectively. Jumping to 100 changes, the time to generate a report increased to 200.2 $\mu$s. Thus at a certain point, additional changelog entries begin to slow down report generation. However as discussed in Section V-B, cookies with large changelog-s could be marked invalid, preventing them from reaching this threshold. Changing the number of Monitored cookies led to greater variability. Generating a report with 10 Monitored cookies took 4 times longer (221.2 $\mu$s) than with a single cookie. With 50 and 100 Monitored cookies, this time balloons to 1,200 and 2,671 $\mu$s respectively. While large, prior works state that websites have between 10 and 20 cookies on average [84, 85, 86]. Therefore we believe these large reports to be unlikely. Further, only Monitored cookies are added to the report. Thus not all of a site’s cookies may be included in the report, minimizing this overhead.
VIII Privacy Considerations
@@
CREAM aims to strike a good balance between security and privacy. Below we discuss privacy-related considerations.
@@
Private/Incognito Browsing. To limit tracking cookies [84, 87, 88], browsers restrict which contexts cookies can be accessed from. While we discuss the Cookie Monster as a single entity, browsers have different Cookie Monsters for different browsing contexts. For example, Incognito browsing has a separate Cookie Monster from standard browsing. Thus, Incognito cookies are not accessible during standard browsing (and vice-versa). As discussed in Section II-C, cookies can also be partitioned to restrict their accessibility by embedding context. As CREAM works on individual cookies, it does not inhibit and remains compatible with these privacy features. Further, CREAM-protected cookies remain manageable by users via the normal browser management interfaces.
@@
Privacy-Enhancing Extensions. Some extensions seek to protect user privacy. Among other checks, they often use the Cookies API to remove known tracking cookies from the browser/requests. As the BrowserOnly attribute hides cookies from extensions, servers could leverage it to avoid privacy-enhancing extensions. As a middle ground, it is possible to allow users to selectively disable the BrowserOnly attribute for specific trusted extensions. This opt-out approach would protect the user by default while allowing discretionary disablement. That said, we note that, regardless of the BrowserOnly attribute, user tracking is still possible through techniques such as browser fingerprinting [89, 90] which are orthogonal to our goal. Further, without approaches to protect cookies, session cookies can be stolen/used, leading to accounts (and all sensitive information therein) being compromised, i.e., a very strong privacy violation. Without browser-based controls such as CREAM, the latter is possible irrespective of privacy-enhancing extensions.
@@
Tracking and Fingerprinting. Monitored cookies rely on symmetric keys to tie a cookie to a browser. However, as keys are generated per site, they cannot be used to track the user across domains. It is also possible to generate keys per origin to limit tracking within a single site, e.g., similar to browsers’ origin isolation [91]. We note that, as session cookies contain the user’s authenticated ID, they identify the user. Thus, irrespective of CREAM, multiple session cookies can track the user across sessions. However, it is possible to update/refresh the keys for each session to prevent their use for cross-session tracking. These keys could also allow individual sites to track users across browsing contexts. This can be prevented by generating keys per site and context rather than per site alone. However, this would incur additional runtime and storage costs.
IX Related Work
@@
Session Protection is a similar external control offered by Flask-Login [95] (Flask’s session management module). When enabled, the server hashes the user’s IP address and user agent (browser) for each session cookie. These hashes are stored on the server and link the cookies to specific hosts. Flask recalculates this hash for all requests with a session cookie and compares it against the saved hash. In the strictest configuration, a hash mismatch invalidates the session. However, by default, a mismatch only marks the session as stale. In this state, the user can still access the site but must re-authenticate for any action that requires a “fresh” session. Flask also supports a separate “Remember Me” token to save and restore sessions. This token is not protected by Session Protection and thus can be stolen to bypass it altogether. As a consequence, the Flask-Paranoid add-on [96] extends Session Protection to the Remember Me token as well. Regardless, while Session Protection can detect cookie theft, it cannot accurately determine if a cookie was stolen as benign actions (i.e., IP address changes) also violate the control. Thus stolen cookies could still provide access to most of a session.
Other proposals use extensions to bolster the security of cookies [4, 97, 98]. CookieExt [4] detects session cookies in incoming requests and marks them as Secure and HttpOnly. Serene [97] maintains a database of session cookie values and ensures the value of outbound cookies are in the database. As discussed in Section I, malicious extensions can bypass existing cookie attributes and reattach cookies to requests. Thus the aforementioned extension-based approaches cannot protect against other extensions.
Many controls change cookies themselves [99, 2, 100, 101]. Signed cookies contain a cryptographic signature in the cookie’s value to support integrity-checking [100] and access-control [101]. Origin cookies [99] add a new “Origin” attribute that greatly limits a cookie’s accessibility to a single origin. One-Time Cookies (OTC) [2] generate a unique token per request rather than per session. OTCs are then tied to a request with a MAC and verified with a shared session key. This session key and any other information needed for verification are stored in a session ticket, encrypted by the server’s secret key. These features limit how cookies are accessed, but do not protect against and still allow theft by malicious extensions. While OTCs reduce the usability of a stolen cookie to a single request, they are enabled through HTTP headers. Thus, malicious extensions can disable OTCs outright and force standard cookie authentication.
@@
X Conclusion
We proposed CREAM: a modification to the web browser’s trust model to protect cookies from malicious extensions. CREAM introduces two new cookie attributes: BrowserOnly and Monitored. BrowserOnly prevents cookies from being accessed by extensions and JS. Monitored maintains accessibility but ties the cookie to a single browser, logs all changes made to the cookie, and sends an authenticated report of changes to the server alongside the cookie. This allows the server to detect and refuse stolen and tampered cookies. We provide an open-source prototype of CREAM built atop Chromium [43] and evaluate it to show that both the BrowserOnly and Monitored attributes incur small overheads.
@@ Appendix A Scraping Browser Extensions
For our analysis of current extensions, we used the Chrome Web Store’s and AMO’s sitemaps [106, 107] to identify potential extension URLs. In total we found 11,490,897 URLs, however, this contained many repeated URLs and subdomains such as an extension’s reviews. Filtering the results, we were left with 172,386 and 533,790 unique Chrome and Firefox extensions respectively. With this, we visited each identified URL and pulled the extension’s name, label, and user count from its page. The majority of extensions were labeled “Themes”. Themes are purely cosmetic and contain no JS or HTML [108]. For this reason, we excluded themes from our analysis. Similarly, 16,380 Chrome extensions were unlabeled. Due to this, we believe these extensions to be Chrome Apps, a now deprecated subclass of extension that behaves similarly to native applications [109]. While Apps are deprecated on most platforms, their deprecation on Enterprise and Education builds of ChromeOS is still ongoing [72]. Since Apps are still technically available we do analyze their permissions, however separately from our main discussion. Firefox also labeled 22 extensions as “Android”. Given this distinction, we also analyze these extensions separately from the main discussion. Filtering out these categories we were left with a dataset of 122,951 Chrome and 37,783 Firefox extensions.
With 177,136 extensions identified (including Chrome Apps and Android variants), the final step was to download each extension’s manifest file. For this, we used an online extension viewer [110] to pull and decompress each extension’s files. Using this tool, we entered each identified extension’s URL and downloaded its manifest file. We repeated this process several times. For each successive pass, we rechecked all URLs that resulted in an error with a larger timeout value. This repetition helped remove false errors due to rate limiting on the server and larger extensions requiring more time to load. When complete, we downloaded 119,152 Chrome and 37,631 Firefox extension manifests. The remaining 3,951 extension URLs resulted in various 400 HTTP errors, suggesting they were either private or had been removed from the repository between our two scrapes. We also pulled 12,351 Chrome App and 22 Android manifests.
TABLE IV: Breakdown of manifest versioning, host permissions, and access to cookies
@@ Appendix B Chrome Apps & Android Analysis
As discussed above, we discovered 16,380 Chrome Apps and 22 Android (Firefox) extensions. From this, we successfully pulled 12,156 App and 22 Android manifests. Similar to our main discussion we analyzed each manifest for its version, host permissions, and access to cookies. A summary of these findings can be found in Table IV.
Nearly all manifests analyzed (99.5%) were written using manifest V2. This is likely due to a lack of incentive to change formats as Firefox currently has no plans to deprecate the format and Chrome Apps will be fully deprecated by the end of 2025 [109]. Looking at host permissions, only 4.3% of manifests declared broad host permissions. However, most Android extensions (81.8%) had “$<$all_urls$>$” permissions. Only 4.1% of Chrome Apps had any broad host permissioning. These limited permissions may be due to Chrome Apps performing more like native applications thus running more locally rather than needing to access many websites.
We saw similar low numbers concerning access to cookies. Only 17 manifests (0.1%) had the Cookies permission. Similarly, only 41 (0.3%) used the WebRequest API with 26 (0.2%) also having access to the WebRequestBlocking permission. While low, 63.6% of Android extensions had the WebRequest permission, and all but one of these also declared the WebRequesBlocking permission. Despite this, manifests with the Cookies, WebRequest, and WebRequestBlocking API affected a combined 2,245,591, 14,699,240, and 13,622,263 users respectively. No manifest declared the DeclarativeNetRequest permission.
We further analyzed which manifests with access cookies also had “$<$all_urls$>$” host permissions. Table V summarizes these findings. Both Android extensions with the Cookies permission could access all cookies in the browser. This was less common in Chrome Apps. Only 20% of Apps with Cookies permissions were able to access all cookies. 92.9% of Android extensions and 48.1% of Chrome Apps with WebRequest permissions could see all traffic in the browser. Similarly, all of these Android extensions and 60% of these Chrome Apps also declare the WebRequestBlocking permission.
TABLE V: Manifests with ”$<$all_urls$>$” and cookie access